OpenStudy (anonymous):

If the graph of y = ax2 + bx + c opens down, which of the following must be true? c > 0 a < 0 c < 0 a > 0

OpenStudy (anonymous):

@whpalmer4 @random231 @TuringTest

OpenStudy (anonymous):

a>0 c>0

OpenStudy (anonymous):

@ravikr.iit its multiple choice

OpenStudy (anonymous):

then only a>0 i thought more than one could be right! :-P

OpenStudy (anonymous):

@ravikr.iit thank you

OpenStudy (anonymous):

@miryvianey :-)

OpenStudy (whpalmer4):

If you have \[y = ax^2+bx+c\] the parabola opens upward if \(a>0\) and downward if \(a<0\) This is easily seen by graphing a few points of \[y = x^2\] and \[y = -x^2\]